# GIFT: Unlocking Full Potential of Labels in Distilled Dataset at Near-zero Cost

### v1

@@ Appendix B Related Work
B.1 Labels in Dataset Distillation
In the field of dataset distillation, the majority of existing methods fix the labels as a one-hot format [33, 42, 40, 3, 32, 14, 18, 41], with a primary focus on optimizing synthetic images. The potential of soft labels has been demonstrated in recent research, and the methods for obtaining soft labels can be broadly categorized into two types.
Optimization-based Soft Labels.
The first type involves learning labels, with several studies [1, 20, 27, 46, 7, 11] finding that making labels learnable can significantly improve performance and efficiency. LD [1] first introduces label distillation and reveals that even only learning labels without learning images can achieve satisfactory performance with high efficiency. Recent work [11] further highlights that optimizing labels can enhance training stability and improve performance.
Teacher model-based soft labels.
The subsequent trajectory of works seek to directly obtain soft labels. Inspired by knowledge distillation [13], TESLA [5]introduces a soft label assignment strategy, directly generating soft labels by leveraging pre-trained teacher models trained on real datasets. These soft labels provide rich intra-class information, thereby improving distillation performance. Following this trend, the state-of-art methods [36, 28, 25] also utilize soft labels predicted by teacher models, achieving significant improvements.
B.2 Knowledge Distillation
A straightforward method to simultaneously utilize soft and hard labels is knowledge distillation [13], which transfers knowledge from a large teacher model to a small student model. In this training process, the student model is supervised by hard labels and soft labels from the teacher’s output. Many following works aim to enhance the use of soft labels for more effective knowledge transfer. [37] investigated the regularization property of soft labels and introduced a teacher-free distillation approach. WSLD [45] analyzes soft labels and distributes different weights for them from a perspective of bias-variance trade-off. DKD [43] decouples the logit and assigns different weights for the target and non-target classes.
Despite the promising potential of knowledge distillation in transferring knowledge from teacher to student models using soft labels, its application to our problem yields limited improvement. A detailed analysis and comparison of these limitations are provided in
@@ Appendix C Experiment Details
We first compare our method GIFT with the state-of-the-art dataset distillation methods, that involve soft labels and can scale to large high-resolution datasets:
  • TESLA [5] marks the first distillation approach that extends to the full ImageNet-1K, circumventing the extensive memory demands associated with MTT-derived methods through a constant memory footprint. However, TESLA does not provide public synthesic datasets, so we are not able to conduct on it.
  • SRe2L [36] and RDED [28]: both of them use soft labels assigned by a teacher model and use them via KL divergence.
  • DATM [11]: initial soft labels assigned by multiple teacher models and then optimized based on trajectory matching. Finally, this method employs soft cross-entropy loss for soft labels.
  • G-VBSM [25]: soft labels are assigned by multiple teacher models and then used via MSE-CE loss function.
Knowledge distillation [13] is a straightforward method to utilize labels, especially for soft labels. Therefore, we also compare our method GIFT with the state-of-the-art knowledge distillation methods that focus on soft labels utilization:
  • KD [13]: it is the first method to transfer knowledge using both hard and soft labels from the teacher’s output.
  • WSLD [45]: it analyzes soft labels and then distributes different weights for them from a perspective of bias-variance trade-off.
  • DKD [43]: it decouples the logits and assigns different weights for the target and non-target classes.
  • NKD [35]: it finds the sum of the two non-target logits is different, preventing logits’ distributions from being identical. Therefore, it normalizes the non-target logits to equalize their sum.
@@
For both dataset distillation and performance evaluation, we employ identical neural network architectures. Consistent with previous studies [3, 5, 44], we use Conv-3 for CIFAR-10 and CIFAR-100 distillation tasks, Conv-4 for Tiny-ImageNet (with the exception of DREAM, which utilizes Conv-3) and ImageNet-1K, Conv-5 for ImageNet-10, and Conv-6 for ImageNet-100 distillation. In line with [3, 5], MTT and TESLA apply a reduced resolution for distilling $224\times 224$ images. According to [36], for retrieving and evaluating distilled datasets, SRe2L and GIFT adopt ResNet-18.
@@
Consistent with recent works [36, 25], the evaluation on the distilled dataset follows the parameters outlined in Table 11. Furthermore, we implement Differentiable Siamese Augmentation (DSA) as described by [39] to enhance images during both the distillation and evaluation phases of our experiments.
@@ Appendix D Experiment Results
Comprehensive Comparison Between Different Loss Functions
Our experiments span two datasets, including Tiny-ImageNet, and ImageNet-1K on $\texttt{IPC}\in\{1,10,50\}$ using ConvNet [11]. Note that the synthetic dataset for DATM on ImageNet-1K is unavailable, precluding comparisons on this dataset. We evaluate at $\texttt{IPC}\in\{1,10,50\}$. The results, visualized in Figure 4, Figure 6, Figure 5 and Figure 7, reveal that the performance of models trained on synthetic datasets is highly sensitive to the choice of the loss function, highlighting the necessity for a unified and effective loss function in dataset distillation.
@@
Ablation study on ImageNet-1K
@@
Appendix E Pytorch Implementation Code

### v2

@@ Appendix B Related Work
B.1 Knowledge Distillation
A straightforward method to simultaneously utilize soft and hard labels is knowledge distillation (Hinton et al., 2015), which transfers knowledge from a large teacher model to a small student model. In this training process, the student model is supervised by hard labels and soft labels from the teacher’s output. Many following works aim to enhance the use of soft labels for more effective knowledge transfer. (Yuan et al., 2020) investigated the regularization property of soft labels and introduced a teacher-free distillation approach. WSLD (Zhou et al., 2021) analyzes soft labels and distributes different weights for them from a perspective of bias-variance trade-off. DKD (Zhao et al., 2022) decouples the logit and assigns different weights for the target and non-target classes.
Despite the promising potential of knowledge distillation in transferring knowledge from teacher to student models using soft labels, its application to our problem yields limited improvement. A detailed analysis and comparison of these limitations are provided in Section 5.3.
@@ Appendix C Experiment Details
To elucidate the rationale behind our method selection for comparison, we categorize current state-of-the-art methods based on two key factors: scalability to ImageNet-1K and the utilization of soft labels. These categorizations are summarized in Table 13.
Table 13: Categorize methods based on their utilization of soft labels and their scalability to ImageNet-1K.
RDED CDA G-VBSM SRe2L DATM SeqMatch DREAM IDC FTD DataDAM MTT DM DSA
Use Soft Label × × × × × × × ×
Scale to ImageNet-1K × × × × × × × ×
Given that our primary focus is on enhancing the use of soft labels in dataset distillation, we restrict our comparisons to methods that involve soft labels:
  • TESLA (Cui et al., 2023) marks the first distillation approach that extends to the full ImageNet-1K, circumventing the extensive memory demands associated with MTT-derived methods through a constant memory footprint. However, TESLA does not provide public synthesic datasets, so we are not able to conduct on it.
  • SRe2L (Yin et al., 2023) and RDED (Sun et al., 2024): both of them use soft labels assigned by a teacher model and use them via KL divergence.
  • DATM (Guo et al., 2024): initial soft labels assigned by multiple teacher models and then optimized based on trajectory matching. Finally, this method employs soft cross-entropy loss for soft labels.
  • G-VBSM (Shao et al., 2024): soft labels are assigned by multiple teacher models and then used via MSE-CE loss function.
  • CDA (Yin & Shen, 2023): soft labels are assigned by a teacher model and are used via soft cross-entropy loss.
Knowledge distillation (Hinton et al., 2015) is a straightforward method to utilize labels, especially for soft labels. Therefore, we also compare our method GIFT with the state-of-the-art knowledge distillation methods that focus on soft labels utilization:
  • KD (Hinton et al., 2015): it is the first method to transfer knowledge using both hard and soft labels from the teacher’s output.
  • WSLD (Zhou et al., 2021): it analyzes soft labels and then distributes different weights for them from a perspective of bias-variance trade-off.
  • DKD (Zhao et al., 2022): it decouples the logits and assigns different weights for the target and non-target classes.
  • NKD (Yang et al., 2023): it finds the sum of the two non-target logits is different, preventing logits’ distributions from being identical. Therefore, it normalizes the non-target logits to equalize their sum.
@@
For both dataset distillation and performance evaluation, we employ identical neural network architectures. Consistent with previous studies (Cazenavette et al., 2022; Cui et al., 2023; Zhao et al., 2023), we use Conv-3 for CIFAR-10 and CIFAR-100 distillation tasks, Conv-4 for Tiny-ImageNet (with the exception of DREAM, which utilizes Conv-3) and ImageNet-1K, Conv-5 for ImageNet-10, and Conv-6 for ImageNet-100 distillation. In line with (Cazenavette et al., 2022; Cui et al., 2023), MTT and TESLA apply a reduced resolution for distilling $224\times 224$ images. According to (Yin et al., 2023), for retrieving and evaluating distilled datasets, SRe2L and GIFT adopt ResNet-18.
@@
Consistent with recent works (Yin et al., 2023; Shao et al., 2024), the evaluation on the distilled dataset follows the parameters outlined in Table 15. Furthermore, we implement Differentiable Siamese Augmentation (DSA) as described by (Zhao & Bilen, 2021) to enhance images during both the distillation and evaluation phases of our experiments.
@@ Appendix D Experiment Results
Application: Continual Learning
Following prior studies (Zhao & Bilen, 2023; Kim et al., 2022; Yin et al., 2023) that leverage synthetic datasets in continual learning to assess the quality of synthetic data, we employ the GDumb framework (Prabhu et al., 2020) for our continual learning setup. This framework sequentially stores prior training data in memory and utilizes both new and stored data for model training.
(a) 5-step
(b) 10-step
Comprehensive Comparison Between Different Loss Functions.
Our experiments span two datasets, including Tiny-ImageNet, and ImageNet-1K on $\texttt{IPC}\in\{1,10,50\}$ using ConvNet (Guo et al., 2024). Note that the synthetic dataset for DATM on ImageNet-1K is unavailable, precluding comparisons on this dataset. We evaluate at $\texttt{IPC}\in\{1,10,50\}$. The results, visualized in Figure 4, Figure 6, Figure 5 and Figure 7, reveal that the performance of models trained on synthetic datasets is highly sensitive to the choice of the loss function, highlighting the necessity for a unified and effective loss function in dataset distillation.
@@
Figure 6: Top-1 accuracy on various synthetic datasets via the SOTA dataset distillation methods across loss functions on Tiny-ImageNet when IPC=50.
@@
(a) CIFAR-100
(b) Tiny-ImageNet
(c) ImageNet-1K
Figure 9: Top-1 accuracy for the SOTA dataset distillation methods on various synthetic datasets with different data augmentation techniques when IPC =10 on ResNet-18 with different $\gamma$.
(a) CIFAR-100
(b) Tiny-ImageNet
(c) ImageNet-1K
Figure 10: Top-1 accuracy for the SOTA dataset distillation methods on various synthetic datasets when IPC =10 on ResNet-18.
Ablation study on ImageNet-1K.
@@
Efficacy of Label Refinement.
The necessity of incorporating hard labels into the soft labels generated by teacher models arises from the inherent limitations in the performance of these teacher models. Specifically, the test accuracies of teacher models are only 61.27%, 49.73%, and 43.6% for CIFAR-100, Tiny-ImageNet, and ImageNet-1K, respectively, when trained on commonly used ConvNet architectures in dataset distillation. Consequently, the accuracy of soft labels is constrained by the suboptimal nature of the teacher models. To mitigate potential inaccuracies in these soft labels, we integrate hard labels to enhance reliability.
To verify the efficacy of the label refinement, we record the label accuracy before and after refinement across each training epoch for three datasets. The parameter $\gamma$, controlling the integration ratio, is set to 0.1. As depicted in Figure 10, refining soft labels results in significant performance improvements of 37.1%, 40.95%, and 71.39% for CIFAR-100, Tiny-ImageNet, and ImageNet-1K, respectively, highlighting the critical role of the proposed label refinement.
@@
SRe2L + Refine 13.3 $\pm$ 0.4 ($\uparrow$ 0.8) 36.3 $\pm$ 0.4 ($\uparrow$ 0.9) 40.5 $\pm$ 0.4 ($\uparrow$ 0.4) 31.8 $\pm$ 0.3 ($\uparrow$ 0.3) 49.7 $\pm$ 0.3 ($\uparrow$ 0.2) 54.5 $\pm$ 0.1 ($\uparrow$ 0.2)
@@
RDED + Refine 21.2 $\pm$ 0.5 ($\uparrow$ 1.2) 38.7 $\pm$ 0.2 ($\uparrow$ 0.2) 42.3 $\pm$ 0.4 ($\uparrow$ 0.5) 42.3 $\pm$ 0.2 ($\uparrow$ 0.9) 55.8 $\pm$ 0.3 ($\uparrow$ 0.3) 59.2 $\pm$ 0.3 ($\uparrow$ 0.4)
@@
G-VBSM + Refine 23.3 $\pm$ 0.4 ($\uparrow$ 0.7) 37.8 $\pm$ 0.3 ($\uparrow$ 0.5) 40.9 $\pm$ 0.3 ($\uparrow$ 0.7) 37.2 $\pm$ 0.3 ($\uparrow$ 0.5) 52.7 $\pm$ 0.2 ($\uparrow$ 0.4) 57.5 $\pm$ 0.2 ($\uparrow$ 0.2)
@@
GIFT Enhance Utilization of Hard and Smoothed Labels
To evaluate the efficacy of GIFT across various data types, we conduct experiments on both distilled and randomly selected datasets, employing hard and smoothed labels, with IPC =10. (1) For the distilled dataset, we use the state-of-the-art dataset distillation method, RDED, which uses soft labels generated by teacher models. In the experiments, we maintain the distilled images and replace soft labels with hard and smoothed labels for model training. (2) For random dataset, we randomly select 10 images for each class from the original dataset. The results for two types of data are presented in Table 18 and Table 19. It is evident that applying our method to label utilization consistently improves performance for both the distilled and the randomly selected data.
Table 18: Evaluation of Loss Functions Across Hard and Smoothed Labels on the Distilled Data Generated via RDED with IPC =10.
Label Type Loss Function CIFAR100 Tiny-ImageNet ImageNet-1K
Hard CE 21.6 $\pm$ 0.2 13.5 $\pm$ 0.3 8.3 $\pm$ 0.4
Ours 22.8 $\pm$ 0.2 ($\uparrow$ 1.2) 14.8 $\pm$ 0.3 ($\uparrow$ 1.3) 9.8 $\pm$ 0.2 ($\uparrow$ 1.5)
Smoothed SoftCE 21.9 $\pm$ 0.2 13.8 $\pm$ 0.2 8.5 $\pm$ 0.2
KL 21.7 $\pm$ 0.3 13.3 $\pm$ 0.3 8.0 $\pm$ 0.3
MSE 22.1 $\pm$ 0.1 14.1 $\pm$ 0.1 8.8 $\pm$ 0.3
Ours 23.1 $\pm$ 0.2 ($\uparrow$ 1.0) 15.2 $\pm$ 0.3 ($\uparrow$ 1.1) 10.2 $\pm$ 0.2 ($\uparrow$ 1.4)
Table 19: Evaluation of Loss Functions Across Hard and Smoothed Labels on the Randomly Selected Data with IPC=10.
Label Type Loss Function CIFAR100 Tiny-ImageNet ImageNet-1K
Hard CE 18.9 $\pm$ 0.3 9.4 $\pm$ 0.1 4.0 $\pm$ 0.3
Ours 20.3 $\pm$ 0.1 ($\uparrow$ 1.4) 11.6 $\pm$ 0.1 ($\uparrow$ 2.2) 5.1 $\pm$ 0.2 ($\uparrow$ 1.1)
Smoothed SoftCE 19.0 $\pm$ 0.3 9.8 $\pm$ 0.3 4.5 $\pm$ 0.2
KL 17.9 $\pm$ 0.2 8.5 $\pm$ 0.3 3.9 $\pm$ 0.1
MSE 19.2 $\pm$ 0.1 9.7 $\pm$ 0.2 4.6 $\pm$ 0.3
Ours 20.4 $\pm$ 0.2 ($\uparrow$ 1.2) 11.8 $\pm$ 0.2 ($\uparrow$ 2.0) 5.6 $\pm$ 0.2 ($\uparrow$ 1.0)
Figure 11: Comparison of Training Loss Between KL and Our GIFT Across Training Epochs
(a) SGD
(b) Adam
(c) AdamW
Figure 12: Gradient Norms Across Various Optimizers and Hyperparameters.
Appendix E Emperical And Theoretical Analysis of Cross-Optimizer Generalization
We begin by analyzing the three optimizers, namely:
  • SGD (Robbins & Monro, 1951) directly computes gradients from loss values, significantly impacting the update of model parameters.
  • Adam (Kingma & Ba, 2014) includes weight decay in the gradient computation, meaning that *when the primary gradient signal (from the loss) is small, weight decay can overshadow it, leading to ineffective updates toward minimizing the loss.
  • AdamW (Loshchilov, 2017) separates the concerns of optimization and regularization. By applying weight decay independently, it ensures that the optimization process remains focused on minimizing the loss function, while regularization acts as a controlled adjustment to the parameter magnitudes.
Optimizers inherently exhibit diverse characteristics, resulting in distinct training dynamics when applied to distilled datasets generated by various methods. Specifically, these distilled datasets, trained with varying loss functions, exhibit distinct loss values. We reveal that the performance of the optimizers is highly influenced by these loss values, as demonstrated by the subsequent empirical evidence and theoretical analysis.
E.1 Emperical Analysis
When employing the KL divergence loss function, the RDED generally exhibits low loss values, as depicted in Figure 11. Notably, our GIFT framework does not achieve small loss values.
To examine the impact of loss values on optimizer performance, we conduct experiments utilizing RDED-generated distilled data. We train models using three distinct optimizers, each with distinct hyperparameter configurations, and compute the gradient norm for each. Specifically, we varied the learning rate for the SGD optimizer and modified the weight_decay parameter for both the Adam and AdamW optimizers.
The gradient norms of these models, presented in Figure 12, demonstrate that when loss values are small, the training dynamics exhibit heightened sensitivity to the optimizer choice. Therefore, the performances of different optimizers are highly influenced by the loss values. Notably, our GIFT framework can not obtain small loss values, achieving robust performance across various optimizers.
E.2 Theoretical Analysis
E.2.1 Stochastic Gradient Descent (SGD)
Stochastic Gradient Descent (SGD) is a foundational optimization algorithm widely used for training machine learning models, particularly neural networks. SGD iteratively updates model parameters to minimize the loss function by moving in the direction of the negative gradient of the loss with respect to the parameters.
SGD Update Rules.
Define the following:
  • $\theta_{t}$: Parameters at time step $t$.
  • $g_{t}=\nabla_{\theta}L(\theta_{t-1})$: Gradient of the loss function $L$ with respect to parameters $\theta$ at time step $t-1$.
  • $\eta$: Learning rate.
The basic SGD update rule is:
$\theta_{t}=\theta_{t-1}-\eta g_{t}$
Sensitivity to Loss Values in SGD.
The gradient $g_{t}=\nabla_{\theta}L(\theta_{t-1})$ indicates the direction and magnitude of change needed to minimize the loss function. Larger loss function values typically result in larger gradients. Therefore, in SGD, both the loss values and the learning rate $\eta$ directly affect the model updates.
E.2.2 Adam Optimizer
Adam (Adaptive Moment Estimation) is an optimization algorithm that combines the advantages of two extensions of stochastic gradient descent: Adaptive Gradient Algorithm (AdaGrad) and Root Mean Square Propagation (RMSProp). Adam maintains per-parameter learning rates adapted based on the first and second moments of the gradients.
Adam Update Rules.
Define the following:
  • $\theta_{t}$: Parameters at time step $t$.
  • $g_{t}=\nabla_{\theta}L(\theta_{t-1})$: Gradient of the loss function $L$ with respect to parameters $\theta$ at time step $t-1$.
  • $m_{t}$: First moment estimate (exponentially decaying average of past gradients).
  • $v_{t}$: Second moment estimate (exponentially decaying average of past squared gradients).
  • $\beta_{1},\beta_{2}$: Decay rates for the first and second moments, respectively.
  • $\epsilon$: Small constant to prevent division by zero.
  • $\eta$: Learning rate.
The update rules are as follows:
$m_{t}$ $=\beta_{1}m_{t-1}+(1-\beta_{1})g_{t}$
$v_{t}$ $=\beta_{2}v_{t-1}+(1-\beta_{2})g_{t}^{2}$
$\hat{m}_{t}$ $=\frac{m_{t}}{1-\beta_{1}^{t}}\quad$ (Bias-corrected first moment)
$\hat{v}_{t}$ $=\frac{v_{t}}{1-\beta_{2}^{t}}\quad$ (Bias-corrected second moment)
$\theta_{t}$ $=\theta_{t-1}-\eta\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$
Weight Decay in Adam.
In the original Adam implementation, weight decay is typically incorporated by adding an $L_{2}$ regularization term directly to the loss function:
$L^{\prime}(\theta)=L(\theta)+\frac{\lambda}{2}\|\theta\|_{2}^{2}$
The gradient of the modified loss function with respect to $\theta$ is:
$\nabla_{\theta}L^{\prime}(\theta)=\nabla_{\theta}L(\theta)+\lambda\theta$
Consequently, the gradient used in the Adam update rule is augmented with the weight decay term:
$g_{t}=\nabla_{\theta}L(\theta_{t-1})+\lambda\theta_{t-1}$
Substituting this into the Adam update equations, the parameter update rule becomes:
$\theta_{t}=\theta_{t-1}-\eta\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$
Here, $\hat{m}_{t}$ and $\hat{v}_{t}$ incorporate the additional gradient component $\lambda\theta_{t-1}$.
E.2.3 AdamW Optimizer
AdamW is a modification of the Adam optimizer that decouples weight decay from the gradient-based update. AdamW addresses the intertwined nature of weight decay and gradient updates in the original Adam, leading to improved generalization performance and more stable training dynamics.
AdamW Update Rules.
The primary distinction in AdamW lies in how weight decay is applied. Instead of incorporating weight decay into the gradient computation, AdamW applies it directly to the parameters after the standard Adam update. The update rule is expressed as:
$\theta_{t}=\theta_{t-1}-\eta\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}-\eta\lambda\theta_{t-1}$
Alternatively, this can be broken down into two sequential steps:
$\theta^{\prime}_{t}$ $=\theta_{t-1}-\eta\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$
$\theta_{t}$ $=\theta^{\prime}_{t}-\eta\lambda\theta_{t-1}$
In this formulation:
  • The first step performs the standard Adam gradient-based update. - The second step applies
  • weight decay independently of the gradient computation.
Sensitivity to Loss Values in AdamW.
By decoupling weight decay from the gradient-based update, AdamW mitigates the issue of weight decay dominating parameter updates when the loss $L(\theta)$ is small. In AdamW, the gradient-based update remains primarily responsible for minimizing the loss, while weight decay independently enforces regularization. This separation ensures that even when $\nabla_{\theta}L(\theta)$ is minimal, the optimizer can continue to adjust parameters based on the loss gradient without being overly constrained by the weight decay term.
When $L(\theta_{t-1})$ is small, $\nabla_{\theta}L(\theta_{t-1})\approx 0$. In AdamW, the update rule is:
$\theta_{t}=\theta_{t-1}-\eta\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}-\eta\lambda\theta_{t-1}$
The gradient-based update $-\eta\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$ remains tied to the loss gradient, allowing continued optimization of $L(\theta)$. Simultaneously, the weight decay term $-\eta\lambda\theta_{t-1}$ independently controls the magnitude of $\theta$ without influencing the direction dictated by the loss gradient. This ensures that weight decay does not overshadow the gradient-based updates, enabling effective model training even when the loss is minimal.
E.2.4 Why Excessive Weight Decay in Adam Impedes Updates When Loss is Small?
Adam’s Update Mechanism Under Small Loss.
Consider the Adam update rule with weight decay integrated into the gradient:
$\theta_{t+1}=\theta_{t}-\eta\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon},$
where $g_{t}=\nabla_{\theta}L(\theta_{t})+\lambda\theta_{t}$ Assume that the loss $L(\theta_{t})$ is sufficiently small, implying $\nabla_{\theta}L(\theta_{t})\approx 0$. Thus $g_{t}\approx\lambda\theta_{t}$. Therefore, the update rule is expressed as:
$\theta_{t+1}\approx\theta_{t}-\eta\frac{\lambda\theta_{t}}{\sqrt{v_{t}}+\epsilon}\approx\theta_{t}\left(1-\frac{\eta\lambda}{\sqrt{v_{t}}+\epsilon}\right),$
Here, the parameter $\theta_{t}$ is scaled by a factor less than 1 (assuming $\eta\lambda/(\sqrt{v_{t}}+\epsilon)>0$), leading to a reduction in $\theta_{t}$. If $\lambda$ is large, the scaling factor can be significantly less than 1, causing $\theta_{t}$ to diminish rapidly. This aggressive shrinking overshadows the limited gradient from the loss function, effectively halting meaningful updates aimed at minimizing $L(\theta)$.
AdamW’s Update Mechanism Under Small Loss.
In AdamW, the update rule is:
$\theta_{t+1}=\theta_{t}-\eta\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}-\eta\lambda\theta_{t}$
Even when $L(\theta_{t})$ is small, the gradient-based update term $-\eta\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$ remains focused on minimizing the loss, while the weight decay term $-\eta\lambda\theta_{t}$ independently enforces regularization.
This separation ensures that:
  • The optimization process remains primarily influenced by the loss gradient.
  • Weight decay controls the magnitude of the parameters without dictating the direction of updates.
Thus, AdamW allows the model to continue optimizing $L(\theta_{t})$ effectively, even when the loss is already minimal, while maintaining controlled regularization through weight decay.
Appendix F Pytorch Implementation Code